7. The futures price of corn is $3.00 and the contract is for 5000 bushels. The purchase margin is $2000, and the maintenance margin is $900. You enter into a contract to deliver corn (at $3.00). The price of corn goes to $3.30. How much money is in your account? Will you have to remit cash because of a margin call? If so, how much?
Answer #1

As one has shorted futures, the margin account becomes=2000+5000*(3-3.30)=500
Yes as margin has fallen below maintenance margin
Amount to be remitted=2000-500=1500